A Bibliography of Publications of Anatol Slissenko

Anatol Slissenko University Paris 12

Informatics 61, Av. du Gen. de Gaulle

CRETEIL, 94010 FRANCE

Tel: 33 (0)1 45 17 16 63 FAX: 33 (0)1 45 17 16 49

E-mail: slissenko@univ-paris12.fr (Internet)

11 October 2020 Version 1.12

Abstract

This bibliography records publications of Anatol Slissenko.

Title word cross-reference

1/ϵ [BGS04]. ^{n} [HKSS93]. R [HKSS93].

12 [MM97]. 12th [MM97].

20th [WH95].

7th [BD97, Win78].

8th [Bec79].

’95 [WH95]. ’97 [BD97]. 98 [Glo98].

AAECC [MM97]. AAECC-12 [MM97]. Abstract [SV08]. Address [Sli78b].

Aleksandrovich [MMM^{+}80, MMOS90]. algebra [MM97]. Algebraic [Glo98, GS97, GS98, GS99, HKSS94, MM97]. Algorithm [GS98, SDM^{+}65, SDM^{+}69, SDM^{+}83a, SDM^{+}83b]. Algorithmic [Sli63, Sli64c, Sli66]. Algorithmical [Sli89]. Algorithms [BS96b, BS97b, BS97a, CS00, EK81, Sli78a, Sli84, MM97]. amidst [GS97, GS98, GS99]. Analysis [Sli89]. Anniversary [MMM^{+}80]. Application [GLS89b]. applications [AM91]. Applied [MM97]. Approach [Sli78a]. Approximating [BGS04]. Approximation [Sli70]. April [AM91, BD97]. Arithmetical [Sli63, Sli64c, Sli67a]. Artificial [OS83, AM91]. August [Glo98, WH95]. Automata [BS98a, BS02]. Automation [SW83].

Based [Sli78b]. Basis [SDM^{+}69, SDM^{+}83b]. Birthday [MMOS90]. borne [HKSS93].

CAAP [BD97]. CAAP/FASE [BD97]. Calculus [SDM^{+}65, SDM^{+}83a]. Can [AVS81]. Certain [Sli67a]. Checking [BS96a, BS96b, BS97b, BS98b, BRS02, SV08]. chemins [HKSS93]. CIM [AM91]. Class [GS97, GS98, GS99]. Classes [BS99]. codes [MM97]. Complexity [BBS95, GLS89c, Sli79, Sli81a, Sli98, BdRS96]. Complexly [Sli73]. Computation [BS98b, Glo98, Sli81a]. Computational [Sli66, Sli79]. Computations [Sli78b]. Computer [GLS89b, SDM^{+}69, SDM^{+}83b, Sli93a, Bec79, EK81, WH95, Win78]. Computing [GS97, GS99]. Concerning [Sli63, Sli64c]. Conference [AM91, BD97]. Considerations [Sli84]. Construction [Sli67b, HKSS93]. Constructive [Sli64b, Sli64a, Sli67b, Sli67c]. Containing [Sli73]. Context [Sli82, Sli92]. Context-Free [Sli82, Sli92]. Continuity [Sli67b, Sli67c, Sli70]. Continuous [BS97a, Sli64a]. correcting [MM97]. Crossing [BS96b, BS97b]. Curriculum [LST85]. Czech [WH95]. Czechoslovakia [Bec79].

dans [HKSS93]. Data [Sli78c]. Decidable [BS98a, BS99, BS02, BRS02]. Decision [BBS95, BdRS96]. Decomposition [Sli92]. Deducibility [SDM^{+}69, SDM^{+}83b]. Deducible [Sli73]. Deduction [SDM^{+}65, SDM^{+}83a]. Describing [Sli82]. Design [GLS89b, GLS89c]. Detection [Sli83]. Determination [SDM^{+}69, SDM^{+}83b]. Development [Sli91b, BD97]. Devising [Sli84]. Diminishing [Sli92]. Discontinuous [Sli64a]. Distributed [CS00]. Do [AVS81]. doubly [BGS04]. Duplexes [Sli63, Sli64c, Sli67a].

Effective [Sli84]. Enumerable [Sli73]. Environment [Sli91b]. error [MM97]. error-correcting [MM97]. Example [Sli64a]. Exhaustive [AVS81]. Expert [GLS89a, GLS89c, Sli91b].

FASE [BD97]. fault [Sli93b]. Finding [HKSS94, Sli80]. Finite [BBS95, Sli78a]. First [BS98a, BS02]. Former [Sli93a]. Formulas [Sli73]. foundations [Bec79, WH95, Win78]. France [BD97, MM97]. Free [Sli82, Sli92]. Functions [Sli67b, Sli67c]. Further [Sli98].

Germany [Glo98]. Grammars [Sli82].

Graph [Sli79].

Hard [Sli82]. High [BS96b, BS97b].

High-Level [BS96b, BS97b]. Homotopy [GS97, GS98, GS99].

Identification [Sli79]. IFIP [AM91]. Increasing [GLS89b]. Industrial [AM91, OS83]. inférieure [HKSS93]. Informatics [LST85]. Information [Sli91a]. Input [Sli76]. Intelligence [OS83, AM91]. Interface [GLS89b]. International [AM91, BD97, Glo98, WH95, MM97]. Inverse [SDM^{+}69, SDM^{+}83b]. ISSAC [Glo98].

Joint [BD97]. June [MM97].

Knowledge [GLS89b, Sli89, Sli91a].

Languages [BS96b, BS97b]. Leningrad [AM91, Sli98]. Level [BS96b, BS97b]. Lille [BD97]. lines [BGS04]. Linguistic [Sli84]. Link [GS97, GS99]. logarithmic [BGS04]. Logic [BS98a, BS98b, BS99, BRS02, Sli98, SV08, BS02]. Logical [SDM^{+}65, SDM^{+}83a, Sli89].

Machine [GLS89b, SDM^{+}65, SDM^{+}83a]. Machines [Sli76, Sli81b, SV08]. Man [GLS89b]. Man-Machine [GLS89b]. Markov [BBS95, BdRS96]. Maslov [DMM^{+}84]. Matching [Sli78c, Sli83]. Mathematical [Bec79, WH95, Win78]. Mathematics [Sli66, EK81]. Maximal [Sli67b, Sli67c, Sli70]. Measures [Sli91a]. Memory [BBS95]. Method [SDM^{+}69, SDM^{+}83b, Sli92]. Metric [Sli64a]. MFCS [WH95]. Minimum [GS97, GS99]. Minimum-Link [GS97, GS99]. Model [BS96a, BS96b, BS97b, BS98b, BRS02, SV08]. Model-Checking [BS96b, BS97b, BRS02, SV08]. Modeling [GLS89a]. Models [Sli78b]. modern [EK81]. Multihead [Sli76].

Natural [SDM^{+}65, SDM^{+}83a]. Nikolai [MMM^{+}80, MMOS90]. Non [Sli64b, Sli64a]. Non-Discontinuous [Sli64a]. Non-Separable [Sli64b].

Obituary [DMM^{+}84]. observed [BdRS96]. Obstacles [GS97, GS98, GS99, HKSS94]. Olomouc [Bec79]. Operations [Sli63, Sli64c, Sli67a]. Operator [Sli64a]. Optimizing [Sli78a]. Order [BS98a, BS02]. Organization [Sli78b].

Palindromes [Sli81b]. partially [BdRS96]. Path [GS97, GS98, GS99, BGS04]. Paths [HKSS94]. Periodicities [Sli80, Sli83]. Petersburg [Sli98]. Petersburg/Leningrad [Sli98]. Plane [GS97, GS98, GS99, HKSS94]. Point [OS83]. Poland [Win78]. Policies [BBS95]. polygonaux [HKSS93]. Polynomial [Sli82]. Polynomial-Time [Sli82]. Polytime [BS96a, BS98b, GS98]. pour [HKSS93]. practice [BD97]. Prague [WH95]. Predicate [BS99, Sli76, SV08]. Probabilistic [BS96a, BS98b]. Probability [BRS02]. Problem [BS96b, BS97b, BS99, Sli78a, Sli89, BGS04]. Problems [AVS81, GLS89c, Sli63, Sli64c, Sli66, Sli81a, Sli82]. Proceedings [Glo98, AM91, Bec79, BD97, EK81, MM97, WH95, Win78]. Processes [BBS95, BdRS96]. Processing [GLS89b, Sli89, Sli91a]. Productivity [GLS89b]. Programming [LST85]. Project [LST85]. Proof [Sli81b]. Properties [Sli78c]. Property [Sli73]. Propositional [SDM^{+}65, SDM^{+}83a]. Proving [Sli78a]. Purpose [GLS89c].

Quality [Sli91a]. Questions [Sli70].

Railroad [BS96b, BS97b]. Real

[Sli78c, Sli80, Sli81b, Sli83]. Real-Time [Sli78c, Sli81b]. reasoning [SW83]. Recognizability [Sli81b]. Recognizing [Sli76]. Reducible [BS98a, BS02]. Refinements [CS00]. Regulators [Sli67b, Sli67c, Sli70]. Related [Sli66]. Representation [Sli89]. Republic [WH95]. Research [Sli93a]. Rostock [Glo98].

Science

[Sli93a, Bec79, EK81, WH95, Win78].

Search [AVS81, SDM^{+}65, SDM^{+}83a, Sli92]. Semantics [BS96b, BS97b, BS97a]. Semi [GS97, GS98, GS99, HKSS94]. Semi-Algebraic [GS97, GS98, GS99, HKSS94]. Separable [Sli64b]. September [Bec79, EK81, WH95, Win78]. Sergei [DMM^{+}84]. Sets [Sli67a, Sli73]. Seventieth [MMOS90]. Shanin [MMM^{+}80, MMOS90]. Shortest [GS98, HKSS94, BGS04]. Simplified [Sli81b]. Simulation [SV08]. Sixtieth [MMM^{+}80]. skew [BGS04]. software [BD97]. Some [Sli63, Sli64c, Sli66, Sli70, Sli78c]. Soviet [Sli93a]. Space [Sli64a]. Spaces [Sli64b]. Special [GLS89c]. Specified [BS98a, BS02]. SSR [EK81]. St [Sli98]. State [SV08]. Storage [Sli78b]. String [Sli78c, Sli79, Sli83]. String-Matching [Sli78c, Sli83]. Structure [Sli78c, Sli89]. Subclasses [Sli82]. Symbolic [Glo98].

Symmetry [Sli76]. Symposium [Bec79, Glo98, WH95, Win78, MM97]. syntax [Sli93b]. System [LST85]. Systems [BS96a, GLS89a, GLS89b, GLS89c, Sli89, Sli91a, Sli91b].

TAPSOFT [BD97]. TC5 [AM91]. TC5/WG5.3 [AM91]. Technological [Sli91b]. Their [BS97b, BS96b]. Theorem [Sli78a]. Theorem-Proving [Sli78a]. Theoretical [Sli93a]. Theory [Sli81a, BD97]. Time [BS97a, BS98a, Sli78c, Sli80, Sli81b, Sli82, Sli83, BS02, BGS04]. Timed [BS96a, BS96b, BS97b, BS98a, BS98b, BS99, CS00, SV08, BS02]. tolerance [Sli93b]. Tool [GLS89b, Sli82]. Toulouse [MM97]. Tree [BS98b]. Turing [Sli76, Sli81b].

Union [Sli93a]. University [Glo98]. Urgench [EK81]. USSR [AM91]. Uzbek [EK81].

Verification [BS98a, BS99, CS00, BS02]. View [OS83, Sli93a].

WG5.3 [AM91]. Word [Sli80].

Years [Sli93a]. Yuryevich [DMM^{+}84].

Zakopane [Win78].

References

Alty:1991:IAA

[AM91] James L. Alty and Leonid I.

Mikulich, editors. Industrial applications of artificial intelligence: proceedings of the IFIP TC5/WG5.3 International Conference on Artificial Intelligence in CIM, Leningrad, USSR, 16–18 April 1990. North-Holland, Amsterdam, The Netherlands, 1991. ISBN 0-444-88981-7. LCCN Q334 .I43 1990.

Adelson-Velski:1981:WCW

[AVS81] G. Adel’son-Vel’ski and A. Slissenko. What can we do with problems of exhaustive search? In Ershov and Knuth [EK81], pages 315–342. CODEN LNCSD9. ISBN 0-387-11157-3 (New York), 3-540-11157-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .A43 1981. URL https:/

/link.springer.com/book/10.
1007/3-540-11157-3. The sym-

posium was organized by the Academy of Sciences of the Uzbek S.S.R.

Beauquier:1995:CFM

[BBS95] D. Beauquier, D. Burago, and A. Slissenko. On the complexity of finite memory policies for Markov decision processes. Lecture Notes in Computer Science, 969:191–200, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1997:TTP

[BD97] M. Bidoit and M. Dauchet, editors. TAPSOFT ’97: theory and practice of software development: 7th International Joint Conference CAAP/FASE, Lille, France, April 14–18, 1997: proceedings, volume 1214 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-62781-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1214.

Burago:1996:CPO

[BdRS96] Dima Burago, Michel de Rougemont, and Anatol Slissenko. On the complexity of partially observed Markov decision processes. Theoretical Computer Science, 157(2):161– 183, May 05, 1996. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.
com/cgi-bin/cas/tree/store/
tcs/cas_sub/browse/browse.
cgi?year=1996&volume=157&issue=
2&aid=2099.

Becvar:1979:MFC

[Bec79] J. Becvar, editor. Mathematical foundations of computer science, 1979: proceedings, 8th Symposium, Olomouc, Czechoslovakia, September 3–7, 1979, volume 74 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNCSD9. ISBN 0-387-09526-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S9194 1979.

Burago:2004:ASP

[BGS04] D. Burago, D. Grigoriev, and A. Slissenko. Approximating shortest path for the skew lines problem in time doubly logarithmic in 1/ϵ. Theoretical Computer Science, 315(2–3):371–404, May 6, 2004. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:2002:LPD

[BRS02] Danièle Beauquier, Alexander Rabinovich, and Anatol Slissenko. A logic of probability with decidable model-checking. Lecture Notes in Computer Science, 2471:306–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer-
                                        ny.com/link/service/series/
                                        0558/bibs/2471/24710306.htm;
                                         http://link.springer-ny.com/
                                        link/service/series/0558/papers/
                                        2471/24710306.pdf.

Beauquier:1996:PMC

[BS96a] D. Beauquier and A. Slissenko. Polytime model checking for timed probabilistic systems. Technical Report 96–08, University Paris 12, Department of Informatics, 1996. 13 pp. URL

                                        http://www.univ-paris12.fr/
                                        lacl/.

Beauquier:1996:RCP

[BS96b] D. Beauquier and A. Slissenko. The railroad crossing problem: Towards semantics of timed algorithms and their model-checking in high-level languages. Technical Report 96–10, University Paris 12, Department of Informatics, 1996. 24 pp. URL http://www.

                                        univ-paris12.fr/lacl/.

Beauquier:1997:SAC

[BS97a] D. Beauquier and A. Slissenko.

On semantics of algorithms with continuous time. Technical Report 96–15 (Revised Version), University Paris 12, Department of Informatics, 1997. 25 pp. URL

http://www.univ-paris12.fr/
lacl/.

Beauquier:1997:RCP

[BS97b] D. Beauquier and A. Slissenko.

The railroad crossing problem: Towards semantics of timed algorithms and their model-checking in high-level languages. Lecture Notes in Computer Science, 1214:201–212, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beauquier:1998:DVR

[BS98a] D. Beauquier and A. Slissenko.

Decidable verification for reducible timed automata specified in a first order logic with time. Technical Report 98–16, University Paris 12, Department of Informatics, 1998. 28 pp. URL

http://www.univ-paris12.fr/
lacl/.

Beauquier:1998:PMC

[BS98b] Danièle Beauquier and Anatol Slissenko. Polytime model checking for timed probabilistic computation tree logic. Acta Informatica, 35(8):645–664, August 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:/

/link.springer-ny.com/link/
service/journals/00236/bibs/
8035008/80350645.htm;   http:
//link.springer-ny.com/link/

                                        service/journals/00236/papers/
                                        8035008/80350645.pdf.

Beauquier:1999:DCV

[BS99] D. Beauquier and A. Slissenko. Decidable classes of the verification problem in a timed predicate logic. In Proc. of the 12th Intern. Symp. on Fundamentals of Computation Theory (FCT’99). Iasi, Rumania, page ???? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., August 30 – September 3 1999. To appear.

Beauquier:2002:DVR

[BS02] Danièle Beauquier and Anatol Slissenko. Decidable verification for reducible timed automata specified in a first order logic with time. Theoretical Computer Science, 275(1–2): 347–388, March 28, 2002. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/gej-ng/10/41/
                                        16/246/27/37/abstract.html.

Cohen:2000:VRT

[CS00] J. Cohen and A. Slissenko. On verification of refinements of timed distributed algorithms. Lecture Notes in Computer Science, 1912:34–??, 2000. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://link.springer-
                                        ny.com/link/service/series/
                                        0558/bibs/1912/19120034.htm;
                                         http://link.springer-ny.com/
                                        link/service/series/0558/papers/
                                        1912/19120034.pdf.

Davydov:1984:SYM

[DMM^{+}84] G. Davydov, Yu. Matiyasevich, G. Mints, V. Orevkov, N. Shanin, and A. Slissenko. Sergei Yuryevich Maslov. obituary. Uspekhi matematicheskikh nauk, 39:239– 240, 1984. CODEN UMANA5. ISSN 0042-1316. In Russian.

Ershov:1981:AMM

[EK81] A. P. Ershov and Donald E.

Knuth, editors. Algorithms in modern mathematics and computer science: proceedings, Urgench, Uzbek SSR, September 16– 22, 1979, volume 122 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNCSD9. ISBN 0-387-11157-3 (New York), 3-540-11157-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA9.58 .A43 1981. URL

https://link.springer.com/
book/10.1007/3-540-11157-3.

The symposium was organized by the Academy of Sciences of the Uzbek S.S.R.

Gloor:1998:IPI

[Glo98] Oliver Gloor, editor. ISSAC 98:

Proceedings of the 1998 International Symposium on Symbolic and Algebraic Computation, August 13–15, 1998, University of Rostock, Germany. ACM Press, New York, NY, USA, 1998. ISBN 1-58113-002-3.

Gorodetsky:1989:ESM

[GLS89a] V. Gorodetsky, A. Lebedev, and A. Slissenko. Expert systems in modeling. In Special Purpose Expert Systems. USSR Ministry of Defense, Moscow, USSR, 1989. Series: Fundamental and Perspective Research in the Interests of the Defense. In Russian.

Gorodetsky:1989:MMI

[GLS89b] V. Gorodetsky, A. Lebedev, and A. Slissenko. Man-machine interface as a tool of increasing the productivity of design and application of computer systems of knowledge processing. In Special Purpose Expert Systems. USSR Ministry of Defense, Moscow, USSR, 1989. Series: Fundamental and Perspective Research in the Interests of the Defense. In Russian.

Gorodetsky:1989:PCD

[GLS89c] V. Gorodetsky, A. Lebedev, and A. Slissenko. Problems of complexity and design of special purpose expert systems. In Special Purpose Expert Systems. USSR Ministry of Defense, Moscow, USSR, 1989. Series: Fundamental and Perspective Research in the Interests of the Defense. In Russian.

Grigoriev:1997:CML

[GS97] D. Grigoriev and A. Slissenko. Computing minimum-link path in a homotopy class amidst semialgebraic obstacles in the plane. Lecture Notes in Computer Science, 1255:114–129, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grigoriev:1998:PAS

[GS98] D. Grigoriev and A. Slissenko.

Polytime algorithm for the shortest path in a homotopy class amidst semi-algebraic obstacles in the plane. In Gloor [Glo98], pages 17–24. ISBN 1-58113-002-

3. URL http://www.acm.org/
pubs/citations/proceedings/
issac/281508/p17-grigoriev/.

Grigoriev:1999:CML

[GS99] D. Grigoriev and A. Slissenko.

Computing minimum-link path in a homotopy class amidst semialgebraic obstacles in the plane. St. Petersburg mathematical journal, 10(2):315–332, 1999. CO-DEN ???? ISSN 1061-0022.

Heintz:1993:BIP

[HKSS93] J. Heintz, T. Krick, A. Slissenko, and P. Solernó. Une borne inférieure pour la construction de chemins polygonaux dans r^{n}. In Publications du département de mathématiques de l’Université de Limoges, pages 94–100. Université de Limoges, Limoges, France, 1993.

Heintz:1994:FSP

[HKSS94] J. Heintz, T. Krick, A. Slissenko, and P. Solernó. Finding shortest paths around semialgebraic obstacles in the plane. Journal of mathematical sciences (New York, N.Y.), 70(4):1944– 1949, 1994. CODEN JMTSEW. ISSN 1072-3374. Russian original in: Zapiski Nauchnykh Seminarov LOMI, 192:164-174, 1991.

Lavrov:1985:CIS

[LST85] S. Lavrov, A. Slissenko, and G. Tseitin. Curriculum for informatics and system programming. project. Mikroprotsessornyye Sredstva i Systemy == Microprocessor Devices and Systems, 4:20–28, 1985. CODEN ???? ISSN 0223-4844. In Russian.

Mora:1997:AAA

[MM97] Teo Mora and H. F. Mattson, editors. Applied algebra, algebraic algorithms, and errorcorrecting codes: 12th international symposium, AAECC-12, Toulouse, France, June 23–27, 1997: proceedings, volume 1255 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-63163-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA268 .A35 1997.

Maslov:1980:NAS

[MMM^{+}80] S. Maslov, Yu. Matiyasevich, G. Mints, V. Orevkov, and A. Slissenko. Nikolai Aleksandrovich Shanin (on his sixtieth anniversary). Uspekhi matematicheskikh nauk, 35(2(212)):241– 245, 1980. CODEN UMANA5. ISSN 0042-1316. In Russian.

Matiyasevich:1990:NAS

[MMOS90] Yu. Matiyasevich, G. Mints, V. Orevkov, and A. Slissenko. Nikolai Aleksandrovich Shanin (on his seventieth birthday). Russian mathematical surveys, 45(1): 239–240, 1990. CODEN ???? ISSN 0036-0279. Russian original in: Uspekhi Matem. Nauk, 45(1):205–206, 1990.

Orlovsky:1983:AII

[OS83] G. Orlovsky and A. Slissenko.

Artificial intelligence: an industrial point of view. In Computers in Design and Manufacturing. Mashinostroyeniye Publ. House, Leningrad, USSR, 1983.

Shanin:1965:AMS

[SDM^{+}65] N. Shanin, G. Davydov, S. Maslov, G. Mints, V. Orevkov, and A. Slissenko. An Algorithm for Machine Search of a Natural Logical Deduction in a Propositional Calculus. Nauka, Leningrad, USSR, 1965. 39 pp.

Shanin:1969:CAD

[SDM^{+}69] N. Shanin, G. Davydov, S. Maslov, G. Mints, V. Orevkov, and A. Slissenko. A computer algorithm for the determination of deducibility on the basis of the inverse method. Zapiski naucnyh seminarov Leningradskogo otdelenija ordena Lenina Matematiceskogo instituta im. V.A. Steklova Akademii nauk SSSR, 16:8–19, 1969. CODEN ZNSLAF. ISSN 0373-2703. In Russian. Translated into English in: The Automation of Reasoning II. Classical Papers on Computational Logic 1967–1970, Springer-Verlag, 1983.

Shanin:1983:AMS

[SDM^{+}83a] N. Shanin, G. Davydov, S. Maslov, G. Mints, V. Orevkov, and A. Slissenko. An algorithm for machine search of a natural logical deduction in a propositional calculus. In Siekmann and Wrightson [SW83], pages 424–483. ISBN 0-387-12043-2 (U.S.: v. 1), 0-387-12044-0 (U.S.: v. 2). LCCN QA76.9.A96 A94 1983. Classical papers on computational logic 1957–1966.

Shanin:1983:CAD

[SDM^{+}83b] N. Shanin, G. Davydov, S. Maslov, G. Mints, V. Orevkov, and A. Slissenko. A computer algorithm for the determination of deducibility on the basis of the inverse method. In Siekmann and Wrightson [SW83], page ??? ISBN 0-387-12043-2 (U.S.: v. 1), 0-387-12044-0 (U.S.: v. 2). LCCN QA76.9.A96 A94 1983. Classical papers on computational logic 1967–1970. Russian original in: Zapiski Nauchnykh Seminarov LOMI, 16:8–19, 1969.

Slissenko:1963:SAP

[Sli63] A. Slissenko. On some algorithmic problems, concerning arithmetical operations on duplexes. Doklady Akademii Nauk SSSR (Soviet Mathematical Doklady), 152(2):292–295, 1963. CODEN DAKNEQ. ISSN 0869-5652. In Russian.

Slissenko:1964:END

[Sli64a] A. Slissenko. An example of nondiscontinuous but not continuous constructive operator in a metric space. Trudy Matematicheskogo instituta imeni V. A. Steklova = Proceedings of the Steklov Institute of Mathematics, 72:524–532, 1964. CODEN TMISAF. ISSN 0081-5438. In Russian.

Slissenko:1964:CNS

[Sli64b] A. Slissenko. On constructive non-separable spaces. Trudy Matematicheskogo instituta imeni V. A. Steklova = Proceedings of the Steklov Institute of Mathematics, 72:533–536, 1964. CO-DEN TMISAF. ISSN 0081-5438. In Russian.

Slissenko:1964:SAP

[Sli64c] A. Slissenko. On some algorithmic problems, concerning arithmetical operations on duplexes. Trudy Matematicheskogo instituta imeni V. A. Steklova = Proceedings of the Steklov Institute of Mathematics, 72:488–523, 1964. CODEN TMISAF. ISSN 0081-5438. In Russian.

Slissenko:1966:SAP

[Sli66] A. Slissenko. Some algorithmic problems related to computational mathematics. In Inter. Congress of Mathematicians, Moscow, 1966. Section 1, Mathematical Logic and Foundations of Mathematics. Abstracts, pages 24–25. ICM, Moscow, USSR, 1966.

Slissenko:1967:AOC

[Sli67a] A. Slissenko. Arithmetical operations on certain sets of duplexes. Trudy Matematicheskogo instituta imeni V. A. Steklova = Proceedings of the Steklov Institute of Mathematics, 93:241– 267, 1967. CODEN TMISAF. ISSN 0081-5438. Russian original in: Trudy Matematicheskogo Instituta Akademii Nauk SSSR, 93:187–207, 1967.

Slissenko:1967:CMC

[Sli67b] A. Slissenko. The construction of maximal continuity regulators for constructive functions. Trudy Matematicheskogo instituta imeni V. A. Steklova = Proceedings of the Steklov Institute of Mathematics, 93:269– 317, 1967. CODEN TMISAF. ISSN 0081-5438. Russian original in: Trudy Matematicheskogo Instituta Akademii Nauk SSSR, 93:208-249, 1967.

Slissenko:1967:MCR

[Sli67c]  A. Slissenko.  On maximal

continuity regulators of constructive functions. Zapiski naucnyh seminarov Leningradskogo otdelenija ordena Lenina Matematiceskogo instituta im. V.A. Steklova Akademii nauk SSSR, 4:201–208, 1967. CODEN ZNSLAF. ISSN 0373-2703. In Russian. Translated into English in: Seminars in Mathematics, V.A.Steklov Mathematical Institute, Leningrad, 4:82–84, 1969.

Slissenko:1970:SQA

[Sli70] A. Slissenko. Some questions of approximation of maximal continuity regulators. Trudy Matematicheskogo instituta imeni V. A.

Steklova = Proceedings of the Steklov Institute of Mathematics, 113:73–78, 1970. CODEN TMISAF. ISSN 0081-5438. In Russian.

Slissenko:1973:PES

[Sli73] A. Slissenko. A property of enumerable sets containing ‘complexly deducible’ formulas. J. of Soviet Mathematics, 1(1):126– 131, 1973. CODEN ???? ISSN ???? Russian original in: Zapiski Nauchnykh Seminarov LOMI, 20:200-207, 1971.

Slissenko:1976:RSP

[Sli76] A. Slissenko. Recognizing a symmetry predicate by multihead Turing machines with input. Trudy Matematicheskogo instituta imeni V. A. Steklova = Proceedings of the Steklov Institute of Mathematics, 129:25– 208, 1976. CODEN TMISAF. ISSN 0081-5438. Russian original in: Trudy Matematicheskogo Instituta Akademii Nauk SSSR, 129:30–202, 1973.

Slissenko:1978:FAP

[Sli78a] A. Slissenko. Finite approach to the problem of optimizing theorem-proving algorithms. J. of Soviet Mathematics, 10(4): 597–603, 1978. CODEN ???? ISSN ???? Russian original in: Zapiski Nauchnykh Seminarov LOMI, 49:123-130, 1975.

Slissenko:1978:MCB

[Sli78b] A. Slissenko. Models of computations based on address organization of storage. In Proc. Soviet Symp. on AI and Automation of Research in Mathematics, Kiev, pages 94–96. Institute of Cybernetics, Kiev, USSR, 1978.

Slissenko:1978:SMR

[Sli78c] A. Slissenko. String-matching in real-time: Some properties of the data structure. In Winkowski [Win78], pages 493–496. CODEN LNCSD9. ISBN 0-387-08917-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S9194 1978.

Slissenko:1979:CCS

[Sli79] A. Slissenko. Computational complexity of string and graph identification. In Becvar [Bec79], pages 182–190. CODEN LNCSD9. ISBN 0-387-09526-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S9194 1979.

Slissenko:1980:FRT

[Sli80] A. Slissenko. Finding in real time of all the periodicities in a word. Soviet Mathematics. Doklady, 21 (2):392–295, 1980. CODEN ???? ISSN 0197-6788. Russian original in: Doklady Akademii Nauk SSSR, 251(1):48–51, 1980.

Slissenko:1981:CPT

[Sli81a] A. Slissenko. Complexity problems of theory of computation. Russian mathematical surveys, 36 (6):23–125, 1981. CODEN ???? ISSN 0036-0279. Russian original in: Uspekhi Matem. Nauk, 36(2):21–103, 1981.

Slissenko:1981:SPR

[Sli81b] A. Slissenko. A simplified proof of real-time recognizability of palindromes on Turing machines. J. of Soviet Mathematics, 15(1): 68–77, 1981. CODEN ???? ISSN ???? Russian original in: Zapiski Nauchnykh Seminarov LOMI, 68:123–139, 1977.

Slissenko:1982:CFG

[Sli82] A. Slissenko. Context-free grammars as a tool for describing polynomial-time subclasses of hard problems. Information Processing Letters, 14(2):52– 56, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Slissenko:1983:DPS

[Sli83] A. Slissenko. Detection of periodicities and string-matching in real time. Journal of Soviet Mathematics, 22(3):1316– 1386, 1983. CODEN JSOMAR. ISSN 0090-4104 (print), 2376-5798 (electronic). Russian original in: Zapiski Nauchnykh Seminarov LOMI, 105:62–173, 1981.

Slissenko:1984:LCD

[Sli84] A. Slissenko. Linguistic considerations in devising effective algorithms. In Proc. Intern. Congress of Mathematicians, August 16– 24, 1983, Waszawa, pages 347– 357. ICM, Waszawa, Poland, 1984.

Slissenko:1989:TPA

[Sli89] A. Slissenko. Towards the problem of analysis of logical and algorithmical structure of the systems of knowledge representation and processing. In Theoretical Aspects and Tools of Applied Intellectual Systems. Scientific Council for Cybernetics of the Acad. Sci. of the USSR, Moscow, USSR, 1989. Series: Questions of Cybernetics. In Russian.

Slissenko:1991:MIQ

[Sli91a] A. Slissenko. On measures of information quality of knowledge processing systems. Information Sciences, 57–58:389–402, 1991. CODEN ISIJBC. ISSN 0020-0255 (print), 1872-6291 (electronic).

Slissenko:1991:TEE

[Sli91b] A. Slissenko. Technological environment for expert systems development. In Alty and Mikulich [AM91], pages 172–175. ISBN 0-444-88981-7. LCCN Q334 .I43 1990.

Slissenko:1992:DSM

[Sli92] A. Slissenko. Diminishing search by the method of context-free decomposition. In R. Yusupov, editor, Methods and Tools of Information Technology in Science and Industry, pages 7– 20. Nauka Publ. House, St.- Petersburg, Russia, 1992. In Russian.

Slissenko:1993:VRY

[Sli93a] A. Slissenko. A view on recent years of research in theoretical computer science in the former Soviet Union. Technique et science informatiques: TSI, 12(1):

9–28, 1993. CODEN TTSIDJ. ISSN 0752-4072, 0264-7419.

Slissenko:1993:FTS

[Sli93b] A. O. Slissenko. On fault tolerance of syntax. Theoretical Computer Science, 119(1):215– 222, October 11, 1993. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.
com/cgi-bin/cas/tree/store/
tcs/cas_sub/browse/browse.
cgi?year=1993&volume=119&issue=
1&aid=1450.

Slissenko:1998:SPL

[Sli98] A. Slissenko. St. petersburg/ leningrad (1961–1998): From logic to complexity and further. In C. Calude, editor, People and Ideas Crafting Theoretical Computer Science, pages 274– 313. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998.

Slissenko:2008:STA

[SV08] A. Slissenko and P. Vasilyev. Simulation of timed abstract state machines with predicate logic model-checking. J.UCS: Journal of Universal Computer Science, 14(12):1984–??, ???? 2008. CODEN ???? ISSN 0948-

6968. URL http://www.jucs.
org/jucs_14_12/simulation_
of_timed_abstract.

Siekmann:1983:AR

[SW83] Jörg H. Siekmann and Graham Wrightson, editors. Automation of reasoning. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-12043-2 (U.S.: v. 1), 0-387-12044-0 (U.S.: v. 2). LCCN QA76.9.A96 A94 1983. Two volumes.

Wiedermann:1995:MFC

[WH95] J. Wiedermann and Petr Hajek, editors. Mathematical foundations of computer science 1995: 20th International Symposium, MFCS ’95, Prague, Czech Republic, August 28–September 1, 1995: proceedings, volume 969 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNCSD9. ISBN 3-540-60246-1 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35 I54 1995.

Winkowski:1978:MFC

[Win78] Jozef Winkowski, editor. Mathematical foundations of computer science, 1978: proceedings, 7th Symposium, Zakopane, Poland, September 4–8, 1978, volume 64 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNCSD9. ISBN 0-387-08917-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .S9194 1978.